# Chain-linked Multiple Matrix Integration via Embedding Alignment

### v2

@@ B.2 Asymmetric matrices integration
    $\mathbf{W}_{\mathbf{X}}^{(i_{\ell-1},i_{\ell})}$ $=\underset{\mathbf{O}\in\mathbb{R}^{d\times d}}{\operatorname{argmin}}\|\hat{\mathbf{X}}^{(i_{\ell-1})}_{\langle\mathcal{U}_{i_{\ell-1}}\cap\mathcal{U}_{i_{\ell}}\rangle}\mathbf{O}-\hat{\mathbf{X}}^{(i_{\ell})}_{\langle\mathcal{U}_{i_{\ell-1}}\cap\mathcal{U}_{i_{\ell}}\rangle}\|_{F},$
    $\mathbf{W}_{\mathbf{Y}}^{(i_{\ell-1},i_{\ell})\top}$ $=\underset{\mathbf{O}\in\mathbb{R}^{d\times d}}{\operatorname{argmin}}\|\hat{\mathbf{Y}}^{(i_{\ell})}_{\langle\mathcal{V}_{i_{\ell-1}}\cap\mathcal{V}_{i_{\ell}}\rangle}\mathbf{O}-\hat{\mathbf{Y}}^{(i_{\ell-1})}_{\langle\mathcal{V}_{i_{\ell-1}}\cap\mathcal{V}_{i_{\ell}}\rangle}\|_{F}.$
@@
[…]
@@
[…]
@@
    Compute $\hat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{V}_{i_{L}}}=\hat{\mathbf{X}}^{(i_{0})}\mathbf{W}^{(i_{0},i_{1})}\mathbf{W}^{(i_{1},i_{2})}\cdots\mathbf{W}^{(i_{L}-1,i_{L})}\hat{\mathbf{Y}}^{(i_{L})\top}$.
0:  $\hat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{V}_{i_{L}}}$.
Algorithm 7 CMMI for overlapping submatrices of an asymmetric matrix
@@
We compare the performance of Algorithm 7 with other matrix completion methods. We simulate a chain of $(L+1)$ overlapping observed submatrices $\{\mathbf{A}^{(i)}\}_{i=0}^{L}$ for the underlying population matrix $\mathbf{P}$ as described in Figure 14, and then predict the yellow unknown block by Algorithm 7. We let all observed submatrices have the same dimension $n\times m$, and let all overlapping parts have the same dimension $(\breve{p}_{n}n)\times(\breve{p}_{m}m)$. For the observed submatrices, we generate the noise matrices $\{\mathbf{N}^{(i)}\}$ by $\mathbf{N}^{(i)}_{st}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\sigma^{2})$ for all $i=0,1,\dots,L$ and all $s\in[n],t\in[m]$, and we let all observed submatrices have the same non-missing probability $q$. For the low-rank underlying population matrix $\mathbf{P}=\mathbf{U}\bm{\Sigma}\mathbf{V}^{\top}$, we randomly generate $\mathbf{U}$ and $\mathbf{V}$ from $\{\mathbf{O}\in\mathbb{R}^{N\times d}\mid\mathbf{O}^{\top}\mathbf{O}=\mathbf{I}\}$ and $\{\mathbf{O}\in\mathbb{R}^{M\times d}\mid\mathbf{O}^{\top}\mathbf{O}=\mathbf{I}\}$, respectively. We fix the rank as $d=3$, and set $\bm{\Sigma}=\text{diag}(N,\tfrac{3}{4}N,\tfrac{1}{2}N)$. We fix the dimensions of the entire matrix at $N\approx 2200$ and $M\approx 2800$, and we vary $L$, the length of the chain, while ensuring that the observed submatrices fully span the diagonal of the matrix Recall that as $L$ increases, we have more observed submarices but each observed submatrix is of smaller dimensions, which then increases the difficulty of recovering the original matrix $\mathbf{P}$. Figure 15 shows the relative $F$-norm estimation error results of recovering the yellow region.
Figure 14: Simulation setting for an asymmetric matrix
@@
Figure 15: The left panel reports relative $F$-norm errors $\|\hat{\mathbf{P}}_{\mathcal{U}_{0},\mathcal{U}_{L}}-\mathbf{P}_{\mathcal{U}_{0},\mathcal{U}_{L}}\|_{F}/\|\mathbf{P}_{\mathcal{U}_{0},\mathcal{U}_{L}}\|_{F}$ for CMMI and other low-rank matrix completion algorithms as $L$ changes. In particular, we vary $L\in\{1,2,3,4,7,9,14,19\}$ while fixing $N\approx 2200$, $M\approx 2800$, $\breve{p}_{n}=\breve{p}_{m}=0.1$, $q=0.8$, $\sigma=0.5$. The results are averaged over $100$ independent Monte Carlo replicates. Note that the averaged relative $F$-norm errors of IRRSV are $\{0.7,8.0,26.1,28.2,26.4,42.9,79.3,211.1\}$ and a portion of these errors are too large to be displayed together with the errors of other algorithms, so we ignore the excessively large part in the left panel. The right panel reports running time for all algorithms and different $L$. For each setting we run $100$ independent Monte Carlo replicates with 10-core parallel computing.
Appendix C Proof of Main Results
C.1 Proof of Theorem 1
@@
We first state two important technical lemmas, one for the error of $\hat{\mathbf{X}}^{(i)}$ as an estimate of the true latent position matrix $\mathbf{X}_{\mathcal{U}_{i}}$ for each $i\in[K]$, and another for the difference between $\mathbf{W}^{(i,j)}$ and $\mathbf{W}^{(i)}\mathbf{W}^{(j)\top}$. The proofs of these lemmas are presented in Section D.2 and Section D.5.
Lemma C.1.
@@
$\mathbf{P}^{(i)}=\mathbf{U}^{(i)}\bm{\Lambda}^{(i)}\mathbf{U}^{(i)\top},\quad\mathbf{A}^{(i)}=\hat{\mathbf{U}}^{(i)}\hat{\bm{\Lambda}}^{(i)}\hat{\mathbf{U}}^{(i)\top}+\hat{\mathbf{U}}^{(i)}_{\perp}\hat{\bm{\Lambda}}^{(i)}_{\perp}\hat{\mathbf{U}}^{(i)\top}_{\perp}.$
Let $\hat{\mathbf{X}}^{(i)}=\hat{\mathbf{U}}^{(i)}(\hat{\bm{\Lambda}}^{(i)})^{1/2}$. Next define $\mathbf{W}^{(i)}$ as a minimizer of $\|\hat{\mathbf{X}}^{(i)}\mathbf{O}-\mathbf{X}^{(i)}\|_{F}$ over all $d\times d$ orthogonal matrices $\mathbf{O}$. Suppose that
@@
    $q_{i}n_{i}\gtrsim\log^{2}{n_{i}},\quad\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}}{q_{i}^{1/2}\lambda_{i,\min}}=\frac{\gamma_{i}}{(q_{i}n_{i})^{1/2}\mu_{i}\log^{1/2}n_{i}}\ll 1.$ (C.1)
@@
$\hat{\mathbf{X}}^{(i)}\mathbf{W}^{(i)}-\mathbf{X}_{\mathcal{U}_{i}}=\mathbf{E}^{(i)}\mathbf{X}_{\mathcal{U}_{i}}(\mathbf{X}_{\mathcal{U}_{i}}^{\top}\mathbf{X}_{\mathcal{U}_{i}})^{-1}+\mathbf{R}^{(i)},$ (C.2)
@@
$\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}\log^{1/2}n_{i}}{q_{i}^{1/2}\lambda_{i,\min}}=\frac{\gamma_{i}}{(q_{i}n_{i})^{1/2}\mu_{i}}\ll 1,$ (C.3)
@@
$\|\hat{\mathbf{X}}^{(i)}\mathbf{W}^{(i)}-\mathbf{X}_{\mathcal{U}_{i}}\|_{2\to\infty}\lesssim\|\mathbf{E}^{(i)}\mathbf{U}^{(i)}\|_{2\to\infty}\cdot\|(\bm{\Lambda}^{(i)})^{-1/2}\|+\|\mathbf{R}^{(i)}\|_{2\to\infty}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})\log^{1/2}n_{i}}{q_{i}^{1/2}\lambda_{i,\min}^{1/2}}=\frac{\gamma_{i}}{(q_{i}n_{i}\mu_{i})^{1/2}}$
@@
Remark 7.
As $\mathbf{P}^{(i)}\in\mathbb{R}^{n_{i}\times n_{i}}$, we generally have $\|\mathbf{P}^{(i)}\|_{F}=\Theta(n_{i})$, e.g., $\mathbf{P}^{(i)}$ has $\Theta(n_{i}^{2})$ entries that are lower bounded by some constant $c_{0}$ not depending on $N$ or $n_{i}$. Thus, as $\mathbf{P}^{(i)}$ is low-rank with bounded condition number, we also have $\lambda_{i,\min}=\Theta(n_{i})$ and the second condition in Eq. (C.1) simplifies to
@@
Similarly, the condition in Eq. (C.3) simplifies to
@@
Both conditions are then trivially satisfied whenever $q_{i}n_{i}\gg\log{n_{i}}$ and $\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i}=O(1)$. Finally when $\lambda_{i,\min}=\Theta(n_{i})$, the bounds in Lemma C.1 simplify to
@@
$\|\hat{\mathbf{X}}^{(i)}\mathbf{W}^{(i)}-\mathbf{X}_{\mathcal{U}_{i}}\|_{2\to\infty}\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})\log^{1/2}n_{i}}{(q_{i}n_{i})^{1/2}}$
@@
with high probability.
We now proceed with the proof of Theorem 1. Recall Eq. (C.2) and let $\xi_{i}:=\hat{\mathbf{X}}^{(i)}\mathbf{W}^{(i)}-\mathbf{X}^{(i)}$ for any $i$. Also denote $\tilde{\mathbf{W}}^{(i,j)}=\mathbf{W}^{(i)\top}\mathbf{W}^{(i,j)}\mathbf{W}^{(j)}$. We then have
$\hat{\mathbf{X}}^{(i)}\mathbf{W}^{(i,j)}\hat{\mathbf{X}}^{(j)\top}-\mathbf{X}_{\mathcal{U}_{i}}\mathbf{X}_{\mathcal{U}_{j}}^{\top}=$ $(\hat{\mathbf{X}}^{(i)}\mathbf{W}^{(i)})(\mathbf{W}^{(i)\top}\mathbf{W}^{(i,j)}\mathbf{W}^{(j)})(\mathbf{W}^{(j)\top}\hat{\mathbf{X}}^{(j)\top})-\mathbf{X}_{\mathcal{U}_{i}}\mathbf{X}_{\mathcal{U}_{j}}^{\top}$
$=$ $(\mathbf{X}_{\mathcal{U}_{i}}+\xi_{i})\tilde{\mathbf{W}}^{(i,j)}(\mathbf{X}_{\mathcal{U}_{j}}+\xi_{j})^{\top}-\mathbf{X}_{\mathcal{U}_{i}}\mathbf{X}_{\mathcal{U}_{j}}^{\top}$
$=$ $(\mathbf{X}_{\mathcal{U}_{i}}+\xi_{i})(\tilde{\mathbf{W}}^{(i,j)}-\mathbf{I})(\mathbf{X}_{\mathcal{U}_{j}}+\xi_{j})^{\top}+(\mathbf{X}_{\mathcal{U}_{i}}+\xi_{i})(\mathbf{X}_{\mathcal{U}_{j}}+\xi_{j})^{\top}-\mathbf{X}_{\mathcal{U}_{i}}\mathbf{X}_{\mathcal{U}_{j}}^{\top}$
$=$ $\xi_{i}\mathbf{X}_{\mathcal{U}_{j}}^{\top}+\mathbf{X}_{\mathcal{U}_{i}}\xi_{j}^{\top}+\xi_{i}\xi_{j}^{\top}+(\mathbf{X}_{\mathcal{U}_{i}}+\xi_{i})(\tilde{\mathbf{W}}^{(i,j)}-\mathbf{I})(\mathbf{X}_{\mathcal{U}_{j}}+\xi_{j})^{\top}$
@@
$\mathbf{R}^{(i,j)}:=\mathbf{R}^{(i)}\mathbf{X}_{\mathcal{U}_{j}}^{\top}+\mathbf{X}_{\mathcal{U}_{i}}\mathbf{R}^{(j)\top}+\xi_{i}\xi_{j}^{\top},\quad\mathbf{S}^{(i,j)}:=(\mathbf{X}_{\mathcal{U}_{i}}+\xi_{i})(\tilde{\mathbf{W}}^{(i,j)}-\mathbf{I})(\mathbf{X}_{\mathcal{U}_{j}}+\xi_{j})^{\top}.$ (C.4)
@@
$\begin{split}&\|\mathbf{R}^{(i,j)}\|_{\max}\leq\|\xi_{i}\|_{2\to\infty}\times\|\xi_{j}\|_{2\to\infty}+\|\mathbf{R}^{(i)}\|_{2\to\infty}\times\|\mathbf{X}_{\mathcal{U}_{j}}\|_{2\to\infty}+\|\mathbf{X}_{\mathcal{U}_{i}}\|_{2\to\infty}\times\|\mathbf{R}^{(j)}\|_{2\to\infty},\\ &\|\mathbf{S}^{(i,j)}\|_{\max}\leq(\|\mathbf{X}_{\mathcal{U}_{i}}\|_{2\to\infty}+\|\xi_{i}\|_{2\to\infty})\times(\|\mathbf{X}_{\mathcal{U}_{j}}\|_{2\to\infty}+\|\xi_{j}\|_{2\to\infty})\times\|\tilde{\mathbf{W}}^{(i,j)}-\mathbf{I}\|.\end{split}$ (C.5)
Next, by Lemma C.1, for any $i$ we have
$\|\mathbf{R}^{(i)}\|_{2\to\infty}\lesssim\frac{\gamma_{i}^{2}}{q_{i}n_{i}\mu_{i}^{3/2}}+\frac{\gamma_{i}}{q_{i}^{1/2}n_{i}\mu_{i}^{1/2}},\quad\|\xi_{i}\|_{2\to\infty}\lesssim\frac{\gamma_{i}}{(q_{i}n_{i}\mu_{i})^{1/2}}$ (C.6)
with high probability. Finally, by Lemma C.2 we have
$\|\tilde{\mathbf{W}}^{(i,j)}-\mathbf{I}\|$ $\lesssim\alpha_{i,j}$ (C.7)
with high probability. Substituting the above bounds in Eq. (C.6) and Eq. (C.7) into Eq. (C.5), we obtain the desired bounds of $\|\mathbf{R}^{(i,j)}\|_{\max}$ and $\|\mathbf{S}^{(i,j)}\|_{\max}$.
C.2 Proof of Theorem 2
Theorem 2 follows the same argument as that for Theorem 1, with the only change being the use of Lemma C.3 below (see Section D.7 for a proof) to bound the difference between $\mathbf{W}^{(i_{0},i_{1})}\cdots\mathbf{W}^{(i_{L-1},i_{L})}$ and $\mathbf{W}^{(i_{0})}\mathbf{W}^{(i_{L})\top}$.
Lemma C.3.
Consider the setting of Theorem 2 and let $\mathbf{T}^{(L)}:=\mathbf{W}^{(i_{0})\top}\mathbf{W}^{(i_{0},i_{1})}\cdots\mathbf{W}^{(i_{L-1},i_{L})}\mathbf{W}^{(i_{L})}$. We then have

### v4

@@ C.2 Asymmetric matrices integration
    $\mathbf{W}_{\mathbf{X}}^{(i_{\ell-1},i_{\ell})}$ $=\underset{\mathbf{O}\in\mathbb{R}^{d\times d}}{\operatorname{argmin}}\|\widehat{\mathbf{X}}^{(i_{\ell-1})}_{\langle\mathcal{U}_{i_{\ell-1}}\cap\mathcal{U}_{i_{\ell}}\rangle}\mathbf{O}-\widehat{\mathbf{X}}^{(i_{\ell})}_{\langle\mathcal{U}_{i_{\ell-1}}\cap\mathcal{U}_{i_{\ell}}\rangle}\|_{F},$
    $\mathbf{W}_{\mathbf{Y}}^{(i_{\ell-1},i_{\ell})\top}$ $=\underset{\mathbf{O}\in\mathbb{R}^{d\times d}}{\operatorname{argmin}}\|\widehat{\mathbf{Y}}^{(i_{\ell})}_{\langle\mathcal{V}_{i_{\ell-1}}\cap\mathcal{V}_{i_{\ell}}\rangle}\mathbf{O}-\widehat{\mathbf{Y}}^{(i_{\ell-1})}_{\langle\mathcal{V}_{i_{\ell-1}}\cap\mathcal{V}_{i_{\ell}}\rangle}\|_{F}.$
@@
    Compute $\begin{aligned} \mathbf{W}^{(i_{\ell-1},i_{\ell})}&=\underset{\mathbf{O}\in\mathbb{R}^{d\times d}}{\operatorname{argmin}}\|\widehat{\mathbf{X}}^{(i_{\ell-1})}_{\langle\mathcal{U}_{i_{\ell-1}}\cap\mathcal{U}_{i_{\ell}}\rangle}\mathbf{O}-\widehat{\mathbf{X}}^{(i_{\ell})}_{\langle\mathcal{U}_{i_{\ell-1}}\cap\mathcal{U}_{i_{\ell}}\rangle}\|_{F}.\end{aligned}$
@@
    Compute $\mathbf{W}^{(i_{\ell-1},i_{\ell})}$ by $\begin{aligned} \mathbf{W}^{(i_{\ell-1},i_{\ell})\top}&=\underset{\mathbf{O}\in\mathbb{R}^{d\times d}}{\operatorname{argmin}}\|\widehat{\mathbf{Y}}^{(i_{\ell})}_{\langle\mathcal{V}_{i_{\ell-1}}\cap\mathcal{V}_{i_{\ell}}\rangle}\mathbf{O}-\widehat{\mathbf{Y}}^{(i_{\ell-1})}_{\langle\mathcal{V}_{i_{\ell-1}}\cap\mathcal{V}_{i_{\ell}}\rangle}\|_{F}.\end{aligned}$
@@
    Compute $\widehat{\mathbf{P}}_{\mathcal{U}_{i_{0}},\mathcal{V}_{i_{L}}}=\widehat{\mathbf{X}}^{(i_{0})}\mathbf{W}^{(i_{0},i_{1})}\mathbf{W}^{(i_{1},i_{2})}\cdots\mathbf{W}^{(i_{L}-1,i_{L})}\widehat{\mathbf{Y}}^{(i_{L})\top}$.
[…]
@@
We compare the performance of Algorithm C.2 with other matrix completion methods. We simulate a chain of $(L+1)$ overlapping observed submatrices $\{\mathbf{A}^{(i)}\}_{i=0}^{L}$ for the underlying population matrix $\mathbf{P}$ as described in Figure C.2, and then predict the yellow unknown block by Algorithm C.2. We let all observed submatrices have the same dimension $n\times m$, and let all overlapping parts have the same dimension $(\breve{p}_{n}n)\times(\breve{p}_{m}m)$. For the observed submatrices, we generate the noise matrices $\{\mathbf{N}^{(i)}\}$ by $\mathbf{N}^{(i)}_{st}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\sigma^{2})$ for all $i=0,1,\dots,L$ and all $s\in[n],t\in[m]$, and we let all observed submatrices have the same non-missing probability $q$. For the low-rank underlying population matrix $\mathbf{P}=\mathbf{U}\bm{\Sigma}\mathbf{V}^{\top}$, we randomly generate $\mathbf{U}$ and $\mathbf{V}$ from $\{\mathbf{O}\in\mathbb{R}^{N\times d}\mid\mathbf{O}^{\top}\mathbf{O}=\mathbf{I}\}$ and $\{\mathbf{O}\in\mathbb{R}^{M\times d}\mid\mathbf{O}^{\top}\mathbf{O}=\mathbf{I}\}$, respectively. We fix the rank as $d=3$, and set $\bm{\Sigma}=\text{diag}(N,\tfrac{3}{4}N,\tfrac{1}{2}N)$. We fix the dimensions of the entire matrix at $N\approx 2200$ and $M\approx 2800$, and we vary $L$, the length of the chain, while ensuring that the observed submatrices fully span the diagonal of the matrix Recall that as $L$ increases, we have more observed submarices but each observed submatrix is of smaller dimensions, which then increases the difficulty of recovering the original matrix $\mathbf{P}$. Figure C.3 shows the relative $F$-norm estimation error results of recovering the yellow region.
Figure C.2: Simulation setting for an asymmetric matrix
@@
Figure C.3: The left panel reports relative $F$-norm errors $\|\widehat{\mathbf{P}}_{\mathcal{U}_{0},\mathcal{V}_{L}}-\mathbf{P}_{\mathcal{U}_{0},\mathcal{V}_{L}}\|_{F}/\|\mathbf{P}_{\mathcal{U}_{0},\mathcal{V}_{L}}\|_{F}$ for CMMI and other low-rank matrix completion algorithms as $L$ changes. In particular, we vary $L\in\{1,2,3,4,7,9,14,19\}$ while fixing $N\approx 2200$, $M\approx 2800$, $\breve{p}_{n}=\breve{p}_{m}=0.1$, $q=0.8$, $\sigma=0.5$. The results are averaged over $100$ independent Monte Carlo replicates. Note that the averaged relative $F$-norm errors of IRRSV are $\{0.9,9.3,26.1,25.0,19.2,34.1,172.0,216.8\}$ and some of these values are too large to be displayed in this panel. The average running time (in log scale) over 100 replicates for algorithms, using 20-core parallel computing and 256 GB memory, is shown in the right panel.
Appendix D Proof of Main Results
D.1 Proof of Theorem 1
@@
We first state two important technical lemmas, one for the error of $\widehat{\mathbf{X}}^{(i)}$ as an estimate of the true latent position matrix $\mathbf{X}_{\mathcal{U}_{i}}$ for each $i\in[K]$, and another for the difference between $\mathbf{W}^{(i,j)}$ and $\mathbf{W}^{(i)}\mathbf{W}^{(j)\top}$. The proofs of these lemmas are presented in Section E.2 and Section E.5.
Lemma D.1.
@@
$\mathbf{P}^{(i)}=\mathbf{U}^{(i)}\bm{\Lambda}^{(i)}\mathbf{U}^{(i)\top},\quad\mathbf{A}^{(i)}=\widehat{\mathbf{U}}^{(i)}\widehat{\bm{\Lambda}}^{(i)}\widehat{\mathbf{U}}^{(i)\top}+\widehat{\mathbf{U}}^{(i)}_{\perp}\widehat{\bm{\Lambda}}^{(i)}_{\perp}\widehat{\mathbf{U}}^{(i)\top}_{\perp}.$
Let $\widehat{\mathbf{X}}^{(i)}=\widehat{\mathbf{U}}^{(i)}(\widehat{\bm{\Lambda}}^{(i)})^{1/2}$, and define $\mathbf{W}^{(i)}=\underset{\mathbf{O}\in\mathcal{O}_{d}}{\arg\min}\|\widehat{\mathbf{X}}^{(i)}\mathbf{O}-\mathbf{X}_{\mathcal{U}_{i}}\|_{F}$. Suppose that
@@
    $q_{i}n_{i}\gtrsim\log^{2}{n_{i}},\quad\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}}{q_{i}^{1/2}\lambda_{i,\min}}=\frac{\gamma_{i}}{(q_{i}n_{i})^{1/2}\mu_{i}\log^{1/2}n_{i}}\ll 1.$ (D.1)
@@
$\widehat{\mathbf{X}}^{(i)}\mathbf{W}^{(i)}-\mathbf{X}_{\mathcal{U}_{i}}=\mathbf{E}^{(i)}\mathbf{X}_{\mathcal{U}_{i}}(\mathbf{X}_{\mathcal{U}_{i}}^{\top}\mathbf{X}_{\mathcal{U}_{i}})^{-1}+\mathbf{R}^{(i)},$ (D.2)
@@
$\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})n_{i}^{1/2}\log^{1/2}n_{i}}{q_{i}^{1/2}\lambda_{i,\min}}=\frac{\gamma_{i}}{(q_{i}n_{i})^{1/2}\mu_{i}}\ll 1,$ (D.3)
@@
$\|\widehat{\mathbf{X}}^{(i)}\mathbf{W}^{(i)}-\mathbf{X}_{\mathcal{U}_{i}}\|_{2\to\infty}\lesssim\|\mathbf{E}^{(i)}\mathbf{U}^{(i)}\|_{2\to\infty}\cdot\|(\bm{\Lambda}^{(i)})^{-1/2}\|+\|\mathbf{R}^{(i)}\|_{2\to\infty}\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})\log^{1/2}n_{i}}{q_{i}^{1/2}\lambda_{i,\min}^{1/2}}=\frac{\gamma_{i}}{(q_{i}n_{i}\mu_{i})^{1/2}}$
@@
Remark D.1.
As $\mathbf{P}^{(i)}\in\mathbb{R}^{n_{i}\times n_{i}}$, we generally have $\|\mathbf{P}^{(i)}\|_{F}=\Theta(n_{i})$, e.g., $\mathbf{P}^{(i)}$ has $\Theta(n_{i}^{2})$ entries that are lower bounded by some constant $c_{0}$ not depending on $N$ or $n_{i}$. Thus, as $\mathbf{P}^{(i)}$ is low-rank with bounded condition number, we also have $\lambda_{i,\min}=\Theta(n_{i})$ and the second condition in Eq. (D.1) simplifies to
@@
Similarly, the condition in Eq. (D.3) simplifies to
@@
Both conditions are then trivially satisfied whenever $q_{i}n_{i}\gg\log{n_{i}}$ and $\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i}=O(1)$. Finally when $\lambda_{i,\min}=\Theta(n_{i})$, the bounds in Lemma D.1 simplify to
@@
$\|\widehat{\mathbf{X}}^{(i)}\mathbf{W}^{(i)}-\mathbf{X}_{\mathcal{U}_{i}}\|_{2\to\infty}\lesssim\frac{(\|\mathbf{P}^{(i)}\|_{\max}+\sigma_{i})\log^{1/2}n_{i}}{(q_{i}n_{i})^{1/2}}$
@@
with high probability, where $\mathbf{W}^{(i)},\mathbf{W}^{(j)}$ are defined in Lemma D.1 and $\mathbf{W}^{(i,j)}=\underset{\mathbf{O}\in\mathcal{O}_{d}}{\operatorname{argmin}}\|\widehat{\mathbf{X}}^{(i)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\mathbf{O}-\widehat{\mathbf{X}}^{(j)}_{\langle\mathcal{U}_{i}\cap\mathcal{U}_{j}\rangle}\|_{F}$.
We now proceed with the proof of Theorem 1. Recall Eq. (D.2) and let $\xi_{i}:=\widehat{\mathbf{X}}^{(i)}\mathbf{W}^{(i)}-\mathbf{X}_{\mathcal{U}_{i}}$ for any $i$. Also denote $\widetilde{\mathbf{W}}^{(i,j)}=\mathbf{W}^{(i)\top}\mathbf{W}^{(i,j)}\mathbf{W}^{(j)}$. We then have
$\widehat{\mathbf{X}}^{(i)}\mathbf{W}^{(i,j)}\widehat{\mathbf{X}}^{(j)\top}-\mathbf{X}_{\mathcal{U}_{i}}\mathbf{X}_{\mathcal{U}_{j}}^{\top}=$ $(\widehat{\mathbf{X}}^{(i)}\mathbf{W}^{(i)})(\mathbf{W}^{(i)\top}\mathbf{W}^{(i,j)}\mathbf{W}^{(j)})(\mathbf{W}^{(j)\top}\widehat{\mathbf{X}}^{(j)\top})-\mathbf{X}_{\mathcal{U}_{i}}\mathbf{X}_{\mathcal{U}_{j}}^{\top}$
$=$ $(\mathbf{X}_{\mathcal{U}_{i}}+\xi_{i})\widetilde{\mathbf{W}}^{(i,j)}(\mathbf{X}_{\mathcal{U}_{j}}+\xi_{j})^{\top}-\mathbf{X}_{\mathcal{U}_{i}}\mathbf{X}_{\mathcal{U}_{j}}^{\top}$
$=$ $(\mathbf{X}_{\mathcal{U}_{i}}+\xi_{i})(\widetilde{\mathbf{W}}^{(i,j)}-\mathbf{I})(\mathbf{X}_{\mathcal{U}_{j}}+\xi_{j})^{\top}+(\mathbf{X}_{\mathcal{U}_{i}}+\xi_{i})(\mathbf{X}_{\mathcal{U}_{j}}+\xi_{j})^{\top}-\mathbf{X}_{\mathcal{U}_{i}}\mathbf{X}_{\mathcal{U}_{j}}^{\top}$
$=$ $\xi_{i}\mathbf{X}_{\mathcal{U}_{j}}^{\top}+\mathbf{X}_{\mathcal{U}_{i}}\xi_{j}^{\top}+\xi_{i}\xi_{j}^{\top}+(\mathbf{X}_{\mathcal{U}_{i}}+\xi_{i})(\widetilde{\mathbf{W}}^{(i,j)}-\mathbf{I})(\mathbf{X}_{\mathcal{U}_{j}}+\xi_{j})^{\top}$
@@
$\mathbf{R}^{(i,j)}=\mathbf{R}^{(i)}\mathbf{X}_{\mathcal{U}_{j}}^{\top}+\mathbf{X}_{\mathcal{U}_{i}}\mathbf{R}^{(j)\top}+\xi_{i}\xi_{j}^{\top},\quad\mathbf{S}^{(i,j)}=(\mathbf{X}_{\mathcal{U}_{i}}+\xi_{i})(\widetilde{\mathbf{W}}^{(i,j)}-\mathbf{I})(\mathbf{X}_{\mathcal{U}_{j}}+\xi_{j})^{\top}.$ (D.4)
@@
$\begin{split}&\|\mathbf{R}^{(i,j)}\|_{\max}\leq\|\xi_{i}\|_{2\to\infty}\times\|\xi_{j}\|_{2\to\infty}+\|\mathbf{R}^{(i)}\|_{2\to\infty}\times\|\mathbf{X}_{\mathcal{U}_{j}}\|_{2\to\infty}+\|\mathbf{X}_{\mathcal{U}_{i}}\|_{2\to\infty}\times\|\mathbf{R}^{(j)}\|_{2\to\infty},\\ &\|\mathbf{S}^{(i,j)}\|_{\max}\leq(\|\mathbf{X}_{\mathcal{U}_{i}}\|_{2\to\infty}+\|\xi_{i}\|_{2\to\infty})\times(\|\mathbf{X}_{\mathcal{U}_{j}}\|_{2\to\infty}+\|\xi_{j}\|_{2\to\infty})\times\|\widetilde{\mathbf{W}}^{(i,j)}-\mathbf{I}\|.\end{split}$ (D.5)
Next, by Lemma D.1, for any $i$ we have
$\|\mathbf{R}^{(i)}\|_{2\to\infty}\lesssim\frac{\gamma_{i}^{2}}{q_{i}n_{i}\mu_{i}^{3/2}}+\frac{\gamma_{i}}{q_{i}^{1/2}n_{i}\mu_{i}^{1/2}},\quad\|\xi_{i}\|_{2\to\infty}\lesssim\frac{\gamma_{i}}{(q_{i}n_{i}\mu_{i})^{1/2}}$ (D.6)
with high probability. Finally, by Lemma D.2 we have
$\|\widetilde{\mathbf{W}}^{(i,j)}-\mathbf{I}\|$ $\lesssim\alpha_{i,j}$ (D.7)
with high probability. Substituting the above bounds in Eq. (D.6) and Eq. (D.7) into Eq. (D.5), we obtain the desired bounds of $\|\mathbf{R}^{(i,j)}\|_{\max}$ and $\|\mathbf{S}^{(i,j)}\|_{\max}$.
D.2 Proof of Theorem 2
Theorem 2 follows the same argument as that for Theorem 1, with the only change being the use of Lemma D.3 below (see Section E.7 for a proof) to bound the difference between $\mathbf{W}^{(i_{0},i_{1})}\cdots\mathbf{W}^{(i_{L-1},i_{L})}$ and $\mathbf{W}^{(i_{0})}\mathbf{W}^{(i_{L})\top}$.
Lemma D.3.
Consider the setting of Theorem 2. Let $\mathbf{T}^{(L)}:=\mathbf{W}^{(i_{0})\top}\mathbf{W}^{(i_{0},i_{1})}\cdots\mathbf{W}^{(i_{L-1},i_{L})}\mathbf{W}^{(i_{L})}$. Then